Evaluate:

11!
___________
(11-6)! 6!

Answer by stanbon(75887)

Evaluate:

11!
___________ = (11!/6!)(1/5!) = 11*10*9*8*7)/(1*2*3*4*5) = 462
(11-6)! 6!
================
Cheers,
Stan H.
====================